OpenStudy (anonymous):

A water balloon is tossed into the air with an upward velocity of 25 ft/s. Its height h(t) in ft after t seconds is given by the function h(t) = − 16t^2 + 25t + 3. After how many seconds will the balloon hit the ground? What will the height be at t = 1 second?

OpenStudy (alekos):

so you need to find t when h(t) = 0

OpenStudy (alekos):

so we have -16t^2 +25t + 3 = 0 now solve for t

OpenStudy (anonymous):

hang on its gonna take me a second

OpenStudy (anonymous):

253/25

OpenStudy (alekos):

no, that's not it

OpenStudy (alekos):

try again

OpenStudy (alekos):

the second part is easy, just find h(1)

OpenStudy (er.mohd.amir):

-16t^2+25+3=0 is a quadratic eq. so use formula for factor it if eq is ax^2+bx+c=0 thhen solution are given x={-b+sqrt (b*b-4ac) }/ 2a and x={-b-sqtr (b*b-4ac)} / 2a